OpenStudy (anonymous):

Find the slope and the y-intercept of the equation 4x + 5y = -9. please help???

OpenStudy (anonymous):

Y=mx+b

OpenStudy (anonymous):

Are you familiar with that formula?

OpenStudy (anonymous):

The slope is -1.600/2.000 The y-intercept is -9/4

OpenStudy (anonymous):

thanks @chasity.ambs

OpenStudy (anonymous):

4x + 5y = -9 5y = -4x - 9 y = -4/5x - 9/5 the slope and the y-intercept of the equation 4x + 5y = -9 are -4/5 and -9/5 respectively. If that is any help @cindycordova1234

jaynator495 (jaynator495):

\(\sf\Huge\rlap{\color{Cyan}{Welcome~To~Open~Study!}}{\color{teal}{\;Welcome~To~Open~Study!}}\) \(\bullet\) Scince You Are New To Openstudy, please Click On This Link And Read The CoC http://openstudy.com/code-of-conduct/ This Will Tell You The Rules Of Openstudy, If You Dont Follow The Rules You Will Get In Trouble \(\bullet\) Scince You Are New To Openstudy, please Click On This Link And Read The TaC http://openstudy.com/terms-and-conditions This Is Legal Stuff... Its Not AS Important But Its Still Important None The Less \(\bullet\) Please Click This Link To Learn LaTeX http://openstudy.com/study#/updates/5491fcc4e4b01cc001531c1d/ With LaTeX you can make Amazing Text Art... You Can Use Hyper Links... If You Think Of It... Most Likely You Can Do It, With This Tutorial Unleash You Inner Creativity And Learn Latex... PS Don't Forget To Medal! ;) \(\bullet\) If You Need Anything Feel Free To Tag Me And I'll Do My Best Help Out

jaynator495 (jaynator495):

Sorry I didn't get here sooner, is there anything you still need help with?

OpenStudy (anonymous):

No problem @cindycordova1234